# Entropy formula for surface diffeomorphisms

### v1

@@ Abstract
Our approach builds on the key ideas developed in the works of Buzzi-Crovisier-Sarig (Invent. Math., 2022) and Burguet (Ann. Henri Poincaré, 2024) concerning the continuity of the Lyapunov exponents.
@@ 1 Introduction
The sequence $\left(\max_{x\in M}\log\left\|Df^{n}_{x}\right\|\right)_{n\geq 1}$ is sub-additive, and hence the above limit exists. For surface diffeomorphisms with large entropy, we establish an equivalent formula for the topological entropy in terms of the volume growth of the tangent cocycle :
@@
  • Note that the integral is taken with respect to the Lebesgue measure. The Lyapunov regular set (that is, the set of points with well-defined Lyapunov exponents) usually has zero Lebesgue measure and therefore does not contribute to the integral above. Even for ergodic volume-preserving Anosov systems-where Lebesgue almost every point has a positive Lyapunov exponent equal to the entropy of the volume measure (Lebesgue measure)-the integral appearing in Theorem A does not converge to the entropy of the volume measure. Instead, it converges to the entropy of the measure of maximal entropy. In this sense, for sufficiently large time $n$, it is the exceptional points that dominate the integral. The order of the operations $\log$ and $\int$ is thus crucial (see Corollary B).
  • In the special case $r=\infty$, no additional term is added to the topological entropy, and our result thus implies the main results of Kozlovski [14] for surface diffeomorphisms. Also note that Theorem A is formulated directly in terms of the derivative $Df^{n}_{x}$, rather than the induced map $(Df^{n}_{x})^{\wedge}$ between exterior algebras of the tangent spaces which are considered by Kozlovski [14]. Although the equivalence of these two quantities for evaluating the topological entropy is not apparent from their respective forms, it is nonetheless natural to work directly with $Df^{n}_{x}$ in the two-dimensional setting, after a more careful analysis. The use of the derivative $Df^{n}_{x}$ in Theorem A might provide a more effective approach for numerically estimating the topological entropy.
@@
    • For $C^{r}$ ($r>1$) surface diffeomorphisms, Burguet [4] also showed that if Lebesgue almost every point has upper Lyapunov exponent strictly larger than $\frac{\lambda^{+}(f)}{r}$, then the system admits an SRB measure.
@@
    • The case of systems with low regularity ($r=1$) is currently not understood by us.
    • The condition $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$ appears to be critical. There are examples for interval maps (see Kozlovski [14]) with $h_{\rm top}(f)<\frac{\lambda^{+}(f)}{r}$, in which the topological entropy vanishes while the volume growth of the tangent cocycle appearing in Theorem A is positive. Although this does not immediately yield counterexamples in our setting of surface diffeomorphisms, we do not expect a fundamental gap between the situations of interval maps and surface diffeomorphisms. We also note that Kozlovski [14, Page 2] mentioned, without providing a precise reference, that Misiurewicz constructed counterexamples for surface diffeomorphisms where the topological entropy is strictly smaller than the volume growth.
@@
The order of the operations $\log$ and $\int$ is crucial in the following sense.
Corollary B.
@@
then there exists finitely many ergodic SRB measures. Moreover, each of them is a measure of maximal entropy.
@@
Again, by Burguet’s result [4], under condition (5), there are countably many ergodic SRB measures $\{\mu_{i}\}_{i\in\mathbb{N}}$ such that Lebesgue almost every $x$ lies in the basin of some $\mu_{i}$ and each $\mu_{i}$ carries maximal entropy. Indeed, in our setting there are only finitely many such ergodic SRB measures, as a direct consequence of the main results of Buzzi-Crovisier-Sarig [7], which assert that there are at most finitely many ergodic measures of maximal entropy.
We also provide a formula expressing the topological entropy in terms of the volume growth of sub-manifolds which extends the classical results in Yomdin theory [27] in $C^{\infty}$ setting.
Theorem C.
We make some comments.
  • In general, beyond our $C^{r}$ setting together with the condition $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$, it is quite subtle to establish an equivalence between the two kinds of volume growth of sub-manifolds in Theorem C :
    This question was implicitly raised in an unpublished manuscript of Buzzi [6, Remark 1.2].
  • If the topological entropy is strictly larger than $\frac{\lambda^{+}(f)}{r}$, then there exists an invariant measure $\mu$ of maximal entropy ([3, 9]). Moreover, by a result of Cogswell [12] (see also Zang [28]), the supremum in Theorem C is attained by the local unstable manifold $W^{u}_{\rm loc}(x)$ for $\mu$-almost every point $x$.
  • It is worth noting that Theorem C actually implies Theorem A via a Fubini-type argument (see the discussions in Section 2.2). However, we choose to prove Theorem A using a more fundamental convex bound, namely Theorem 2.2. This approach offers a deeper insight into the mechanism governing the volume growth of tangent cocycle. There is no essential difference between these two approaches, since both ultimately follow from Theorem 2.1.
We believe that the ideas of Theorem A and Theorem C might extend to other topological invariants (see Llibre–Saghin [17] for a survey of various invariants measuring dynamical complexity), such as the growth rate of hyperbolic periodic points. This has been established by Burguet [2] in the $C^{\infty}$ setting and it was conjectured by Burguet in the $C^{r}$ case:
Conjecture ([2]).
@@
For any $C^{r}$ ($r>1$) diffeomorphism $f$ on a compact surface $M$ with $h_{\rm top}(f)>\frac{\lambda^{+}(f)}{r}$. For any $0<\delta<h_{\rm top}(f)$,
@@
2 Convex bounds for volume growth
In this section, we present the main results used to prove Theorem A and Theorem C.
2.1 Mechanism behind convex bounds for volume growth

### v2

@@ Abstract
We also characterize the topological entropy via the volume growth of curves and several applications are presented. Our approach builds on the key ideas developed in the works of Buzzi-Crovisier-Sarig (Invent. Math., 2022) and Burguet (Ann. Henri Poincaré, 2024) concerning the continuity of the Lyapunov exponents.
MSC Class : 37A05; 37B40; 37C40; 37D25; 37E30
@@ 1.1 Statement of the main results
The sequence $\left\{\max_{x\in M}\log\left\|Df^{n}_{x}\right\|\right\}_{n\geq 1}$ is sub-additive, and hence the above limit exists.
For surface diffeomorphisms with large entropy, we establish an equivalent formula for the topological entropy in terms of the volume growth of the tangent cocycle :
@@
  • Note that the integral is taken with respect to the Lebesgue measure. The Lyapunov regular set (that is, the set of points with well-defined Lyapunov exponents) usually has zero Lebesgue measure and therefore does not directly contribute to the integral above. Even for ergodic volume-preserving Anosov systems-where Lebesgue almost every point has a positive Lyapunov exponent equal to the entropy of the volume measure (Lebesgue measure)-the integral appearing in Theorem A does not converge to the entropy of the volume measure. Instead, it converges to the topological entropy. In this sense, for sufficiently large time $n$, it is the exceptional points (those for which the derivative has not yet grown at the rate of the positive Lyapunov exponent up to time $n$) that dominate the integral. The order of the operations $\log$ and $\int$ is crucial (see Corollary C).
  • In the special case $r=\infty$, no additional term is added to the topological entropy, and our result thus implies the main results of Kozlovski [23] for surface diffeomorphisms. Also note that Theorem A is formulated directly in terms of the derivative $Df^{n}_{x}$, rather than the induced map $(Df^{n}_{x})^{\wedge}$ between exterior algebras of the tangent spaces which is considered by Kozlovski [23]. Although the equivalence of these two quantities for evaluating the topological entropy is not apparent from their respective forms, it is nonetheless natural to work directly with $Df^{n}_{x}$ in the two-dimensional setting, after a more careful analysis. The use of the derivative $Df^{n}_{x}$ in Theorem A might provide a more effective approach for numerically estimating the topological entropy.
    The extension of the formula in Theorem A from the $C^{\infty}$ to the $C^{r}$ setting is far from routine and requires significant improvements in Yomdin theory. The key obstruction lies in the presence of local complexity (indicated by the Yomdin term $\frac{\lambda^{+}(f)}{r}$) . In the $C^{\infty}$ case, this obstruction disappears which drastically simplifies the problem  [^1] This is also the case for Anosov systems, where no local complexity is present, and the formula in Theorem A can be proved in a relatively straightforward way, without any assumption on the topological entropy or the regularity (see, e.g., Kifer-Newhouse [21], Yang-Zang [34]).. In contrast, in the $C^{r}$ setting, the persistence of a nontrivial local complexity term makes the analysis substantially more delicate. We refer to the background in Section 1.2 for a detailed comparison between the two settings.
@@
    • For $C^{r}$ ($r>1$) surface diffeomorphisms, Burguet [5] also showed that if Lebesgue almost every point has maximal Lyapunov exponent strictly larger than $\frac{\lambda^{+}(f)}{r}$, then the system admits an SRB measure.
@@
    • The case of systems with low regularity ($r=1$) is not understood by us.
    • The condition $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$ appears to be critical. There are examples for interval maps (see Kozlovski [23]) with $h_{\rm top}(f)<\frac{\lambda^{+}(f)}{r}$, in which the topological entropy vanishes while the volume growth of the tangent cocycle appearing in Theorem A is positive. Although this does not immediately yield counterexamples in our setting of surface diffeomorphisms, we do not expect a fundamental gap between the situations of interval maps and surface diffeomorphisms  [^2] We also note that Kozlovski [23, Page 2] mentioned, without providing a precise reference, that Misiurewicz constructed counterexamples for surface diffeomorphisms where the topological entropy is strictly smaller than the volume growth..
@@
We also provide a formula expressing the topological entropy in terms of the volume growth of the sub-manifolds which extends the classical results in Yomdin theory [35] in $C^{\infty}$ setting.
Theorem B.
We make some comments.
  • In general, beyond our $C^{r}$ setting together with the condition $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$, it is quite subtle to establish an equivalence between the two kinds of volume growth of the sub-manifolds in Theorem B :
    This question was implicitly raised in an unpublished manuscript of Buzzi [7, Remark 1.2].
  • If the topological entropy is strictly larger than $\frac{\lambda^{+}(f)}{r}$, then there exists an invariant measure $\mu$ of maximal entropy ([4, 10]). Moreover, by a result of Cogswell [13] (see also Zang [36]), the supremum in Theorem B is attained by the local unstable manifold $W^{u}_{\rm loc}(x)$ for $\mu$-almost every point $x$. In fact, Corollary E shows that the supremum is attained for any ergodic measure $\mu$ with $h(f,\mu)>\frac{\lambda^{+}(f)}{r}$, and not only for measures of maximal entropy as one might expect.
  • It is worth noting that Theorem B actually implies Theorem A. Indeed, by a Fubini-type argument (see the proof of Theorem 2.2 in Section 5.1), there exists a constant $K>0$ such that for any $n\geq 1$,
    As a consequence, by Theorem B, we have
    $\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)=h_{\rm top}(f).$
    However, we will directly prove Theorem A using a more fundamental maximum bound, namely Theorem 2.2. This approach offers a deeper insight into the mechanism governing the volume growth of tangent cocycle. But there is no essential difference between these two approaches, since both ultimately follow from Theorem 2.1.
1.2 Background and applications
@@
Next, we present two applications (Corollary C, Corollary D) of Theorem A, as well as two applications (Corollary E, Corollary F) of Theorem B.
The order of the operations $\log$ and $\int$ in Theorem A is crucial in the following sense.
Corollary C.
@@
then there exists finitely many ergodic SRB measures. Moreover, each of them is a measure of maximal entropy  [^7] A measure of maximal entropy is an $f$-invariant probability measure whose metric entropy is equal to the topological entropy $h_{\rm top}(f)$..
@@
Again, by Burguet’s result [5], under condition (5), there are at most countably many ergodic SRB measures $\{\mu_{i}\}_{i\in\mathbb{N}}$ such that Lebesgue almost every $x$ lies in the basin of some $\mu_{i}$ and each $\mu_{i}$ carries maximal entropy. Indeed, in our setting there are only finitely many such ergodic SRB measures, as a direct consequence of the main results of Buzzi-Crovisier-Sarig [8], which assert that there are at most finitely many ergodic measures of maximal entropy.
The following result, while not new  [^9] Although this result does not appear in this exact form, Corollary D can also be deduced from Burguet’s result [5] by considering SRB measures, in a way similar to the proof of Corollary C. Here we give a technically lighter proof using Theorem A., is another interesting consequence of Theorem A. Denote the usual maximal Lyapunov exponent of $x$ by
$\lambda^{+}(f,x):=\limsup_{n\to+\infty}\frac{1}{n}\log\left\|Df^{n}_{x}\right\|.$
Corollary D.
For any $C^{r}$ ($r>1$) diffeomorphism $f$ on a compact surface $M$ with $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$,
$\lambda^{+}(f,x)\leq h_{\rm top}(f)$
for Lebesgue almost every point $x$.
Proof.
By a general result in measure theory (see Lemma 6.5) and Theorem A,
$\lambda^{+}(f,x)\leq\lim_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx=h_{\rm top}(f)$
for Lebesgue almost every point $x$. ∎
The following Corollay E is an application of Theorem B, showing that the maximal volume growth is attained w.r.t. any ergodic measure $\mu$ satisfying $h(f,\mu)>\frac{\lambda^{+}(f)}{r}$, rather than being restricted to measures of maximal entropy. This conclusion is somewhat unexpected. For instance, consider an SRB measure $\mu$: although its unstable manifolds carry $\mu$-typical points with full induced Lebesgue measure, their volume growth does not converge to the positive Lyapunov exponent (which coincides with the entropy of $\mu$), but instead to the topological entropy.
It is instructive to compare this phenomenon with the so-called Almost Anosov systems constructed by Hu-Young [17], where Lebesgue almost every point has zero Lyapunov exponents, yet a measure of maximal entropy still exists. In particular, a curve along which Lebesgue almost every point has zero Lyapunov exponents may nevertheless exhibit maximal volume growth, equal to the topological entropy. On the other hand, it is not hard to see: a curve along which Lebesgue almost every point has positive Lyapunov exponents cannot have zero volume growth.
The proof of Corollary E relies on background from homoclinic relations (see [8]) and Pesin theory (see [1]). Here we only provide an outline of the argument, omitting the technical details.
Corollary E.
@@
For any $C^{r}$ ($r>1$) transitive diffeomorphism $f$ on a compact surface $M$ and any ergodic measure $\mu$ with $h(f,\mu)>\frac{\lambda^{+}(f)}{r}$, we have for $\mu$-almost every $x$,
$h_{\rm top}(f)=\lim_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(x)\right)\right)$
where $W^{u}_{\rm loc}(x)$ denotes the local unstable manifold at $x$.
Outline of the proof.
Since $f$ is transitive, by [8, Theorem 6.12], there is a unique homoclinic class supports all ergodic measures $\mu$ with $h(f,\mu)>\frac{\lambda^{+}(f)}{r}$. Moreover, this homoclinic class supports a unique ergodic measure $\nu$ of maximal entropy  [^10] The existence of measures of maximal entropy for the $C^{r}$ case is established by Burguet [4, Corollary 1] and the uniqueness is established by Buzzi-Crovisier-Sarig [8, Main Theorem Revisited].. Hence $\mu$ must be homoclinically related to $\nu$, i.e., there are two subsets $R_{\mu},R_{\nu}$ with $\mu(R_{\mu}),\nu(R_{\nu})>0$ such that for any $x\in R_{\mu},y\in R_{\nu}$, the unstable manifold $W^{u}(x)$ intersects the stable manifold $W^{s}(y)$ transversely at some points, and likewise $W^{s}(x)$ intersects $W^{u}(y)$ transversely at some points. Note that for $\nu$-almost every $y\in R_{\nu}$,
$\nu_{y}\left(R_{\nu}\right)>0$
where $\nu_{y}$ denotes the conditional measure  [^11] In fact, conditional measures are defined with respect to subordinate partitions and we disregard this point here. See Ledrappier-Young [18] for further details. of $\nu$ on $W^{u}_{\rm loc}(y)$. We fix any such $y\in R_{\nu}$. Let $z$ be a transverse point of $W^{u}(x)$ and $W^{s}(y)$. Let $m_{1}\in\mathbb{N}$ be large enough such that
$z\in f^{m_{1}}\left(W^{u}_{\rm loc}\left(f^{-m_{1}}\left(x\right)\right)\right).$
By Inclination Lemma  [^12] The Inclination Lemma [8, Lemma 2.7] requires $y$ to be recurrent. As $\nu$-almost every point in $R_{\nu}$ is recurrent (see [8, Section 2.3(e),(f)]), we may assume without loss of generality that $y$ is recurrent., there are $m_{2}\in\mathbb{N}$ and some curve $D$ with
$z\in f^{-m_{2}}\left(D\right)\subset f^{m_{1}}\left(W^{u}_{\rm loc}\left(f^{-m_{1}}\left(x\right)\right)\right)$
such that $D$ is sufficiently $C^{1}$-close to $W^{u}_{\rm loc}(y)$. Consequently, the stable holonomy map $\pi_{s}:\Lambda\to D$ is well defined on some Pesin set  [^13] Here $\Lambda$ refers to the standard Pesin set, where points admit stable and unstable manifolds with uniform size, and the angle between the stable and unstable directions is uniformly bounded away from zero. See [1] for details on Pesin theory. $\Lambda\subset W^{u}_{\rm loc}(y)$ with $\nu_{y}(\Lambda)>0$. See Figure 1.
Figure 1: The stable holonomy map $\pi_{s}$
Let us first briefly recall the argument in [36] (see also [13], [27] for similar ideas) establishing
$h(f,\nu)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(y)\right)\right).$ (6)
The fact that $\nu_{y}(\Lambda)>0$ is what ensures sufficient volume growth. Given any small $\varepsilon>0$ and any large $n\in\mathbb{N}$, since $\nu_{y}\left(\Lambda\right)>0$, by a fiber-version of Katok’s entropy formula (see [36, Theorem A (1)]), there is a maximal $(n,\varepsilon)$-separated subset $\Omega_{n,\varepsilon}$ of $\Lambda$ with
$\Omega_{n,\varepsilon}\gtrsim e^{n\cdot\left(h(f,\nu)-\varepsilon\right)}$
such that for any $w\in\Omega_{n,\varepsilon}$,
${\rm Vol}\left(f^{n}\left(B^{u}(w,n,\varepsilon)\right)\right)\gtrsim e^{-n\varepsilon}$
where $B^{u}(w,n,\varepsilon)$ denotes the unstable dynamical ball,
$B^{u}(w,n,\varepsilon):=\left\{w^{\prime}\in W^{u}_{\rm loc}(w):\ d\left(f^{k}(w^{\prime}),f^{k}(w)\right)<\varepsilon\ \text{ for all }0\leq k<n\right\}.$
As a consequence,
${\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(y)\right)\right)\gtrsim\#\Omega_{n,\varepsilon}\cdot\min_{w\in\Omega_{n,\varepsilon}}{\rm Vol}\left(f^{n}\left(B^{u}(w,n,\varepsilon)\right)\right)\gtrsim e^{n\cdot\left(h(f,\nu)-2\varepsilon\right)}$
which proves (6) by the arbitrariness of $\varepsilon$. The key observation is that the same argument applies with $D$ in place of $W^{u}_{\rm loc}(y)$. Indeed, since $D$ is connected to $W^{u}_{\rm loc}(y)$ through the stable holonomy on $\Lambda$, it follows that, for any large $n$ and any $w\in\Lambda$, the point $f^{n}\left(\pi_{s}(w)\right)$ remains inside the Pesin chart centered at $f^{n}(w)$. As a consequence,
${\rm Vol}\left(f^{n}\left(B^{u}\left(\pi_{s}(w),n,\varepsilon\right)\right)\right)\approx{\rm Vol}\left(f^{n}\left(B^{u}(w,n,\varepsilon)\right)\right)\gtrsim e^{-n\varepsilon}.$
Hence
${\rm Vol}\left(f^{n}\left(D\right)\right)\gtrsim\#\Omega_{n,\varepsilon}\cdot\min_{w\in\Omega_{n,\varepsilon}}{\rm Vol}\left(f^{n}\left(B^{u}\left(\pi_{s}(w),n,\varepsilon\right)\right)\right)\gtrsim e^{n\cdot\left(h(f,\nu)-2\varepsilon\right)}.$
This shows for any $x\in R_{\mu}$,
$h_{\rm top}(f)=h(f,\nu)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(D\right)\right)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}\left(f^{-m_{1}}\left(x\right)\right)\right)\right).$
By Theorem B  [^14] Any local unstable manifold is contained in a curve from the family in Theorem B, as local unstable manifolds are nearly straight at small scales.,
$h_{\rm top}(f)=\lim_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}\left(f^{-m_{1}}\left(x\right)\right)\right)\right).$
By ergodicity and the backward contraction of the local unstable manifolds, this equality above in fact holds for $\mu$-almost every $x$ without passing to backward iterates (see Lemma 6.6):
$h_{\rm top}(f)=\lim_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(x)\right)\right).$
The proof of Corollary E in fact yields the following more general statement: if two hyperbolic ergodic measures $\mu$ and $\nu$ are homoclinically related, then for $\mu$-almost every point $x$,
$h(f,\nu)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}\left(x\right)\right)\right).$
As a consequence of variational principle, for any hyperbolic ergodic measure $\mu$ supported on a homoclinic class $H(\mathcal{O})$ and for $\mu$-almost every $x$,
$h_{\rm top}\left(f,H(\mathcal{O})\right)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(x)\right)\right)$
where $h_{\rm top}\left(f,H(\mathcal{O})\right)$ denotes the topological entropy of the homoclinic class.
In general, the topological entropy of a curve is bounded above by its volume growth (see Lemma 6.2), and the inequality can be strict. Indeed, topological entropy only reflects the number of dynamical balls, capturing the global complexity, whereas volume growth depends not only on this number but also on the geometric distortion within each dynamical ball, such as folding and bending, which represent local complexity. However, the following corollary shows that when a curve carries full topological entropy, its volume growth attains the maximal rate which coincides with the topological entropy. In this case, the contribution of local complexity inside dynamical balls becomes negligible.
Write $\log^{+}x:=\max\{\log x,\,0\}$.
Corollary F.
[…]
$h_{\rm top}(f)=\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right).$
Proof.
By Lemma 6.2, we have
$h_{\rm top}(f)=h_{\rm top}\left(f,\sigma\right)\leq\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right).$
On the other hand, by Theorem B, the right-hand side is bounded above by $h_{\rm top}(f)$. Therefore, equality holds. ∎
1.3 Some further questions
We believe that the ideas of Theorem A and Theorem B might extend to other topological invariants (see Llibre-Saghin [24] for a survey of various invariants measuring dynamical complexity), such as the growth rate of hyperbolic periodic points. This has been established by Burguet [3] in the $C^{\infty}$ setting and it was conjectured by Burguet in the $C^{r}$ case:
Question 1 ([3]).
For any $C^{r}$ ($r>1$) diffeomorphism $f$ on a compact surface $M$ with $h_{\rm top}(f)>\frac{\lambda^{+}(f)}{r}$ and any $0<\delta<h_{\rm top}(f)$,
@@ 1.3 Some further questions
We also propose some questions related to the formula in Theorem A:
$h_{\rm top}(f)=\lim_{n\to+\infty}a_{n}:=\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx.$
Question 2.
In general, does $a_{n}$ always converge ? If so, what is its limit if it is not the topological entropy ?
Remark.
Note that $a_{n}$ might NOT be a sub-additive sequence, even when $f$ preserves the volume (i.e., when the Lebesgue measure is an invariant measure).
We believe that a deeper formula underlies Theorem A.
Question 3.
Is it reasonable to ask about the rate of convergence in Theorem A ?
Remark.
Note that the definition of $a_{n}$ already contains the factor $\frac{1}{n}\log$”. It might therefore be too ambitious to expect a specific rate of convergence (for instance, exponential). However, we do not know this convincingly. We also note that Buzzi-Crovisier-Sarig [10] have recently shown that measures of maximal entropy are exponentially mixing.
Denote by ${\rm Diff}^{r}(M)$ the space of all $C^{r}$ diffeomorphisms on $M$ equipped with the $C^{r}$ topology. Define the functions $\varphi_{n},\varphi:{\rm Diff}^{r}(M)\to\mathbb{R}$:
$\varphi_{n}\left(g\right):=\frac{1}{n}\log\int_{M}\left\|Dg^{n}_{x}\right\|\,dx,\qquad\varphi:=\limsup_{n\to+\infty}\varphi_{n}.$
Question 4.
Is $\varphi$ upper semi-continuous ?
Remark.
[…]
Question 5.
Does an analogue of Theorem A hold for interval maps ?
Remark.
This is known to hold for $C^{\infty}$ interval maps by the work of Kozlovski [23]. It remains unclear whether new obstructions arise when extending it to $C^{r}$ interval maps. In this paper, we only consider diffeomorphisms.
It is also worth noting that there have been several recent works on the continuity of Lyapunov exponents and the existence of SRB measures for interval maps (see, for example, Delplanque [14], Delplanque-Li [15], and Li [19]).
2 Maximum bounds for volume growth
In this section, we present the main results used to prove Theorem A and Theorem B.
2.1 Mechanism behind the maximum bounds for volume growth